A balloon is inflated in at 325 K to a volume of 2 liter. When it is taken into the walk-in refrigerator where the temperature is 275K, what will the new volume be?
1
Expert's answer
2019-10-22T05:43:41-0400

Assuming pressure to be constant.

So,"\frac{V_1}{T_1}=\frac{V_2}{T_2}" ;

"V_1=2L;T_1=325 K;T_2=275 K;" Substituting the values,

"\frac{2}{325}=\frac{V_2}{275}\implies V_2=\frac{275\times 2}{325}=1.69L"